OpenStudy (jmansedg):

Matt and Emily are building a rectangular vegetable garden for their grandmother. They can use up to 60 feet of fencing to enclose the garden. Which of the following is a possible combination of length and width they can use to build the garden? A. Length: 5 feet, Width: 26 feet B. Length: 11 feet, Width: 20 feet C. Length: 12 feet, Width: 17 feet D. Length: 17 feet, Width: 14 feet

OpenStudy (jmansedg):

wiil fan if right

OpenStudy (jmansedg):

@TheCuteOne25

OpenStudy (thecuteone25):

I think B but i'm not 100% sure

OpenStudy (jmansedg):

the next one is Determine which of the following ordered pairs is the solution of the given system. x + 3y = 10 2x - y = 6 A. (4, 2) B. (7, 1) C. (16, -2) D. (3, 0)

OpenStudy (jmansedg):

help me please

OpenStudy (isaidavila):

I think it's A

OpenStudy (jmansedg):

ok what about this one Select the system below which meets the following conditions: The slope-intercept forms are identical, so the graphs are identical. Both equations share all their ordered pairs as solutions. A. y + 3x = 5 y = 2x + 5 B. y = 4x - 1 y = 3x - 2 C. y = 2x - 4 y/2 + 2 = x D. None of the above

OpenStudy (jmansedg):

mate hurry i have to turn this in

OpenStudy (isaidavila):

x+3y=10 3(2x-y)=18 You multiply by 3 to eliminate y and find x Then sum up the equations and 7x=28 x=4

OpenStudy (jmansedg):

and i still have 5 more

OpenStudy (isaidavila):

It's the C bro just solve for y the seco d equation